If (log a) / (b-c) = (log b) / (c-a) = (log c) / (a-b) then a^b+c.b^c+a.c^a+b is equal to
 

Grade:11

1 Answers

Mayukh Chakrabarty
6 years ago
Assume base to be 10 and each ratio equals k. Then a = 10^k(b-c) hence a^(b+c) = 10^k(b²-c²) . Similarly b^(c+a) = 10^k(c²-a²) and c^(a+b) = 10^k(a²-b²)so answer is 10^k(b²-c²+c²-a²+a²-b²) = 1